OpenStudy (anonymous):

What is the integral of (secx)/(9tan^2(x))

zepdrix (zepdrix):

\[\large \int\limits \frac{\sec x}{9\tan^2x}dx\]Hmm there doesn't appear to be a nice `U substitution` so let's convert everything to sines and consines.

OpenStudy (anonymous):

u-substitution doesnt work

zepdrix (zepdrix):

\[\large \int\limits\limits \frac{\left(\dfrac{1}{\cos x}\right)}{9\left(\dfrac{\sin^2x}{\cos^2x}\right)}dx \qquad = \qquad \frac{1}{9}\int\limits\limits \left(\frac{1}{\cos x}\right)\left(\frac{\cos^2x}{\sin^2x}\right)dx\]

zepdrix (zepdrix):

Understand how to proceed gal? :) Do some cancellations with a couple cosines then you can do a nice easy `u sub` from there i think :D

OpenStudy (anonymous):

simplify from zepdrix's last equation gives you \[1/9\int\limits_{}^{}(cotx * cscx)dx\]

OpenStudy (anonymous):

which is simply -1/9 * cscx + c

OpenStudy (anonymous):

Hmm that's what I got but Webassign says its wrong.

OpenStudy (anonymous):

what's the website

zepdrix (zepdrix):

Hmm well the answer is correct. Must just be a formatting error. Are you typing otu the csc? Or using the button for it?

OpenStudy (anonymous):

parenthesis aruond x.. write -1/9 * csc(x) + c